A Gallup poll to survey the top concerns of Americans was conducted. Suppose that 743 women and 713 men were independently and randomly selected, and that 383 women and 282282 men chose the state of the economy as their biggest concern. Can we conclude that the proportion of women ( p1 ), choosing the state of the economy as their biggest concern, exceeds the proportion of men ( p2 )?  Use a significance level of α=0.1 for the test.

Step 1 of 6 : State the null and alternative hypotheses for the test.

Step 2 of 6: Find the values of the two sample proportions, pˆ1 and pˆ2. Round your answers to three decimal places.

Step 3 of 6: Compute the weighted estimate of p, p‾ . Round your answer to three decimal places.

Step 4 of 6: Compute the value of the test statistic. Round your answer to two decimal places.

Step 5 of 6: Find the P-value for the hypothesis test. Round your answer to four decimal places.

Step 6 of 6: Make the decision to reject or fail to reject the null hypothesis
Answer #1

positive & please give a 12 = 713 D-743 t = 383, x2= 282 Ho - b P7R Hi 516 7 / hi 396 X2 P = x+x2 383+282 7431213 hith -0.4579.594 p value p vakre a 2 (10) Null hypothesis will be rejected Null Hypo. Will be rejected